OpenStudy (anonymous):

I need help understanding the set-up to this problem dealing with vectors: A 100-lb weight hangs from two wires. First wire (T1) forms an angle with the ceiling of 50 degrees, second wire (T2) forms an angle with the ceiling of 32 degrees. Find the tensions T1 and T2 in both wires and the magnitudes of the tensions.

OpenStudy (anonymous):

The example is from the stewart text - don't need the answer just help understanding the set-up. It goes on explaining that the tensions can first be written mathematically as \[\left| T1 \right|= -\left| T1\right|\cos(50) i + \left| T1 \right|\sin(50)j\] \[\left| T2 \right|= \left| T2\right|\cos(32) i + \left| T2 \right|\sin(32)j\] Any help refreshing my memory on why cos and sin appear in these equations would be much appreciated!

OpenStudy (aum):